OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x)

OpenStudy (tkhunny):

Root of 'a' implies factor of (x-a). Go!

OpenStudy (anonymous):

could you explain more?

OpenStudy (tkhunny):

Root of 3 implies factor of (x-3).

OpenStudy (anonymous):

so would it only be (x+2)(x-3)

OpenStudy (tkhunny):

Excellent, excepting the word "multiplicity", What shall we do with that?

OpenStudy (anonymous):

multiply both sides by 2 ?

OpenStudy (tkhunny):

Good guess, but you've only one side. Just repeat yourself. Root of 3 implies factor of (x-3) Root of 3 multiplicity 2 implies factor of (x-3) twice, either (x-3)(x-3) or more simply \((x-3)^{2}\)

OpenStudy (anonymous):

ooh okok once i get the exponent on both sides what do i do ?? so i just multiply -3 *-3?

OpenStudy (tkhunny):

Where are you getting the "both sides"? We are creating an algebraic expression. There's only one side. Let's try this. Give these zeros, create a polynomial with these zeros. -2 3 (x+2)(x-3) Give these zeros, create a polynomial with these zeros. -2 3 3 (x+2)(x-3)(x-3) = (x+2)(x-3)^2 Give these zeros, create a polynomial with these zeros. -2 3 (multiplicity 2) (x+2)(x-3)(x-3) = (x+2)(x-3)^2 Give these zeros, create a polynomial with these zeros. -2 3 (multiplicity 5) (x+2)(x-3)^5